If the adjacent is 9cm and I want to find the hypotenuse,with an angle of 60 degrees, what method do I use. I've tried cos(9/60) and the same with negative cos but it doesn't work.
1
Expert's answer
2011-12-06T09:36:22-0500
If the adjacent is 9cm and an angle is 60 degrees, then hypotenuse is
C = 9/cos(60) = 9/0.5 = 18cm.